OpenStudy (ciel21):

*will give medal if you help with all questions please~~* (questions will be in the comments)

OpenStudy (ciel21):

simplify 73 - 2 × (3 - 1)

OpenStudy (anonymous):

73 -2 (2) 73-4 69

OpenStudy (ciel21):

301 - 4[(3 + 1)^3 - (3^3 + 7)]

OpenStudy (anonymous):

The problem to solve is: 301 - 4[(3 + 1)^3 - (3^3 + 7)] 3+1 evaluates to 4 Raising 4 to the exponent 3 is the same as multiplying 4 by itself 3 times. Multiply 4 by 4 Multiply 4 and 4 1 4 × 4 = 16 16 Multiply 16 by 4 Multiply 16 and 4 1 16 × 4 = 64 64 (3+1)^3 evaluates to 64 3^3 evaluates to 27 3^3+7 evaluates to 34 (3+1)^3-(3^3+7) evaluates to 30 Multiply 4 and 30 1 4*[(3+1)^3-(3^3+7)] evaluates to 120 301-4*[(3+1)^3-(3^3+7)] evaluates to 181 The final answer is 181

OpenStudy (ciel21):

The temperature of a city at sunset was -1°F. Overnight, the temperature decreased by 9°F. What was the lowest temperature overnight in that city?

OpenStudy (anonymous):

Decreased so subtract Therefore, -1-9=-10<----Final answer

OpenStudy (anonymous):

-10 F

OpenStudy (ciel21):

Griffin ordered a pair of sneakers online. He had a $19 credit that he applied toward the purchase, and then he used a credit card to pay for the rest of the cost. If the shoes cost $73, how much did Griffin charge to his credit card when he bought the sneakers?